OpenStudy (jaredstone4):

Let f(x) = sqrt(x) If the rate of change of f at x=c is twice the rate of change at x=1, then c= A) 1/4 B) 1 C) 4 D) 1/sqrt(2) E) 1/2sqrt(2)

OpenStudy (radar):

Have you studied derivatives yet?

OpenStudy (jaredstone4):

Yes, a few chapters ago. This is on a review packet for our final and I just forget how to do it

OpenStudy (radar):

f(x) = sqrt(x)

OpenStudy (radar):

f'(x) =(1/2)x^-(1/2)

OpenStudy (radar):

that is the rate of change or the derivative. Now substitute the value of 1 for x solve and then calculate twice to answer the problem.

OpenStudy (radar):

I used the "power rule" to obtain the derivative.............remember how that went? does this refresh your memory? ax^b the derivative would be abx^(b-1)

OpenStudy (jaredstone4):

Yes the power rule I remember. I remember most of what we learned but I often have trouble just setting up the problem

OpenStudy (jaredstone4):

Okay so I got f'(1) = 1/2. So, since f'(x) is twice f'(1), the answer is 1. Correct?

OpenStudy (radar):

For some unknown reason, I can't use the Equation or the Draw buttons beneath the box Yes you are on the correct path.

OpenStudy (radar):

sqrt(x)=x^(1/2) the derivative would be 1/2x^-(1/2) which can be shown as 1/(2 sqrt x)

OpenStudy (jaredstone4):

Yes I got that much..then subbing in I got f'(1) = 1/2. 2(f'(1)) = c so c =1. Is that the final answer?

OpenStudy (radar):

I don't see any errors, good luck with your studies.

OpenStudy (jaredstone4):

Thank you very much for your help!

OpenStudy (radar):

you're welcome